OpenStudy (momo_loves_cte):

Aneel has 4 64-ounce cartons of orange juice and 2 48-ounce cans of apple juice. How many ounces of juice does he have altogether? 256 ounces 304 ounces 352 ounces 382 ounces

OpenStudy (love_ranaa):

To find the answer do 64*4=? and add 48*2=?

OpenStudy (love_ranaa):

64 x 4= 256 48 x 2= 96 256+96 = 352